OpenStudy (anonymous):

Look at Craig's budget below. Find the percentages associated with each category of his budget: Monthly net income: $1,300 Expenses: Housing: $325.00 Transportation: $150.00 Utilities: $60.00 Food: $400.00 Savings: $100.00 Other Expenses: $265.00

OpenStudy (anonymous):

plz need help

OpenStudy (texaschic101):

monthly net income = 1300 housing = 325/1300 = 0.25 = 25% transportation = 150/1300 = 0.115 = 11.5% utilities = 60/1300 = 0.046 = 4.6% do you understand how I am doing these ?

OpenStudy (anonymous):

no i still dont get can you plz explain

OpenStudy (texaschic101):

part/whole....and the whole amount, in this problem, is always gonna be 1300 so for housing....put the housing amount and divide it by the whole amount 325/1300....now divide them....use a calculator and turn this into a decimal 325/1300 = 0.25.....now turn it into a percent by moving the decimal two spaces to the right...or you can just multiply it by 100....thus, giving you 25% does that help clear things up ?

OpenStudy (texaschic101):

for food, it would be 400/1300....then turn to a decimal, then a percent

OpenStudy (anonymous):

some what so im dividing the net monthly income with each expense than turning the decimal into a percent correct

OpenStudy (texaschic101):

correct...now your problem does not say round percent to nearest whole number....so, I am guessing you do not have to round them.

OpenStudy (texaschic101):

any more problem with this and you can just tag me :)

OpenStudy (anonymous):

for food i got 0.307692 if i change it it should 30.78 @texaschic101

OpenStudy (anonymous):

right